A 1.0 kg object is brought to planet Mercury where the acceleration due to gravity is 0.38 times its value on Earth. (a) What is the weight of the object on Earth? (b) What is the mass of the object on Mercury? What is its weight on Mercury?




1
Expert's answer
2022-05-30T15:18:21-0400

(a)

"W_E=mg_E\\
=\rm 1.0\: kg\times 9.8\: N/kg=9.8\: N"

(b)

"W_M=mg_M\\
=\rm 1.0\: kg\times 0.38\times 9.8\: N/kg=3.7\: N"

Mass of the object is the same on both planets.